Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Fri Nov 7 11:08:57 MET 1997) Success
No known problems

Onsala (A. Polatidis at Fri Nov 7 11:16:06 MET 1997) Success
No known problems. We are recording only LCP so signal only at ODD video converters. Tape will be sent to Bonn correlator by DHL on Friday 7/11..

Effelsberg (dgraham at Fri Nov 7 11:40:59 MET 1997) Minor failures
preob was missing due to -1 preob time in drg. put tsys in postob from 1943 on.

Noto (C.Trigilio at Fri Nov 7 12:59:27 MET 1997) Minor failures
Upconverter out of order; substitute with a racaldana. No problem during the observations. Clear sky.

Torun (K.Borkowski at Fri Nov 7 13:03:42 MET 1997) Success

Westerbork (tony foley at Fri Nov 7 14:11:59 MET 1997) Minor failures
Looks OK. No phasecal signal.

Jodrell (Alastair Gunn at Sat Nov 8 17:42:13 MET 1997) Success
FS crashed during source change and restarted at 17:45. Track 21 is dead. No other known problems. Tape HSL00555 sent by DHL on Friday 07/11/97.

MPIfR (dgraham at Mon Nov 10 12:54:00 MET 1997) Minor failures
correlator report -good fringe on Eff-Noto ,waiting for other tapes

MPIfR (dgraham at Mon Nov 10 18:16:31 MET 1997) Minor failures
corelator report Monday 18MEZ- amplitudes in units of 0.01percent correlation are - BN 67 BJ 88 BS 33 , cannot find W, L no tape yet.

MPIfR (dgraham at Tue Nov 11 12:15:58 MET 1997) Minor failures
correlator report 3. Tapes from Medicina and Torun arrived, correlation OK. Torun track 7 (vlba 10) has bits but no signal. Delays found for all stations match GPS data.

Urumqi (Li Suqin at Fri Nov 14 10:39:43 MET 1997) Success
No problem. It is fine day. Hmaser to GPS=2.95us, Format to GPS=2.70us.

Calibration (Fredrik T Rantakyro at Tue Dec 9 14:55:20 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name c97c4a.antab Read the comments in the file to get further information.

JIVE (Lorant Sjouwerman at Wed Dec 17 16:54:47 MET 1997) Success
Network monitoring report of C-band session (first experiment)